OpenStudy (anonymous):

Simplify the given expression. 1/4(6+20y)-1/2(19-8y)

OpenStudy (anonymous):

9y - 8

OpenStudy (anonymous):

?

OpenStudy (anonymous):

(6 + 20y - 2[19 - 8y])/4 = (6 + 20y - 38 + 16y)/4 = (36y - 32)/4 = 9y - 8. Answer:

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }(6+20y)-\frac{ 1 }{ 2 }(19-8y)=\frac{ 1 }{ 4 }((6+20y)-2(19-8y) )=\frac{ 1 }{ 4 }(44+36y)\] =(11+4y)

OpenStudy (anonymous):

(11+9y) sorry